# Stick to your Role! Stability of Personal Values Expressed in Large Language Models

### v1

@@ Abstract
The standard way to study Large Language Models (LLMs) through benchmarks or psychology questionnaires is to provide many different queries from similar minimal contexts (e.g. multiple choice questions). However, due to LLM’s highly context-dependent nature, conclusions from such minimal-context evaluations may be little informative about the model’s behavior in deployment (where it will be exposed to many new contexts). We argue that context-dependence should be studied as another dimension of LLM comparison alongside others such as cognitive abilities, knowledge, or model size. In this paper, we present a case-study about the stability of value expression over different contexts (simulated conversations on different topics), and as measured using a standard psychology questionnaire (PVQ) and a behavioral downstream task. We consider 19 open-sourced LLMs from five families. Reusing methods from psychology, we study Rank-order stability on the population (interpersonal) level, and Ipsative stability on the individual (intrapersonal) level. We explore two settings: with and without instructing LLMs to simulate particular personalities. We observe similar trends in the stability of models and model families - Mixtral, Mistral and Qwen families being more stable than LLaMa-2 and Phi - over those two settings, two different simulated populations, and even in the downstream behavioral task. When instructed to simulate particular personas, LLMs exhibit low Rank-Order stability, and this stability further diminishes with conversation length. This highlights the need for future research directions on LLMs that can coherently simulate a diversity of personas, as well as how context-dependence can be studied in more thorough and efficient ways. This paper provides a foundational step in that direction, and, to our knowledge, it is the first study of value stability in LLMs. The project website with code is available at https://sites.google.com/view/llmvaluestability.
Keywords: Large Language Models; LLM Stability; Context-dependence; Personal Values; Simulated populations
@@ 1 Introduction
In recent years, there has been an emergence of research using psychological tools to study Large Language Models (LLMs). In those studies, LLMs have often been used to simulate populations by instructing them to simulate different personas (?, ?). LLMs have also been used without providing such instructions, i.e. treated as a participant in a human study (?, ?). Questions in such studies have revolved around how Language Models express values (?, ?), personality traits (?, ?, ?), cognitive abilities (?, ?), and how they could replicate human data (?, ?).
@@
Figure 1: How do LLM’s expressed values change as a function of context? An LLM is first prompted to play a specific role (e.g. Gandalf). Then, a conversation on a topic (e.g. joke) with an interlocutor model (same LLM prompted to simulate a human user) is generated. Then, the LLM simulating the persona is given a psychology questionnaire aimed to assess its expressed values. We study the stability of these expressed values across diverse conversation topics and lengths. We consider various personas to be simulated, as well as the case when the LLM is not prompted to play any particular persona. The messages and instructions in gray are set manually, and the messages in white are generated.
@@
The use of psychological tools with LLMs opens up many scientific questions, for example regarding the nature of how the text generated by LLMs depends on context, i.e. information present in prompts or prior interaction with the model. Prompts or prior interaction, which we denote here as ’context’, can include any textual information such as instructions (e.g. personas to simulate), the dialogue history, stories written in specific styles, etc. Such contexts guide the generation of text by LLMs: different contexts may result in the expression of different behavior and values. This is sometimes beneficial and expected, e.g. an instruction to simulate some persona should influence the behavior and expressed values to be more aligned with that persona. However, this can also be problematic, e.g. a specific conversation topic may drastically influence the expressed behavior and values in unexpected ways as we will highlight in experiments below. It should be noted that, depending on the application, different types of context-dependence can be beneficial or not.
@@
The issue of unexpected context dependence in LLM is of crucial importance. Standard evaluation benchmarks, including those using psychological questionnaires to assess properties of LLMs, consist of sets of queries, often presented with a similar minimal context (e.g. knowledge or value-related questions presented as multiple choice questions with limited context). When deployed, LLMs are exposed to many new unforeseen contexts. This means that the standard benchmarks, by themselves, cannot estimate a model’s behavior in deployment (due to the LLMs’ highly context dependent nature). It is therefore crucial to evaluate the robustness of different models to unexpected context-based changes in behavior.
@@
This challenge is particularly acute with the use of psychological questionnaires aimed at measuring psychological dimensions like values. Those tools were initially designed to probe humans, and thus make various assumptions about humans: for example, it is expected that the answers of most humans to questionnaires about value preferences should not be significantly influenced by the content of a randomly picked Wikipedia article shown to them beforehand. As we will show below, such an assumption does not hold for many LLMs, and thus strongly limits general interpretability of using these questionnaires in a context-independent manner. It is thus key to understand better how LLMs’ behavior (e.g. expression of values) may maintain coherence or change as a function of various kinds of contexts (ranging from explicit instruction to play a particular persona to discussions about topics that seem unrelated to the expressed psychological dimensions one studies).
@@
Previous research included certain experiments regarding unwanted context-based change (usually regarding syntactic changes in the prompt). These experiments led to conflicting results, sometimes showing robustness (?, ?, ?, ?, ?), and sometimes sensitivity (?, ?, ?). These inconclusive results motivate research into the nature and the extent of context-dependence of various LLMs.
@@
In this paper, we present a case-study focusing on studying the stability of value expressed in 19 open-source LLMs from 5 families. First, we study to what extent can LLMs simulate various personas in coherent ways, i.e. expressed values should change according to the instructed persona, but not based on the topic of a conversation not related to values. In other words, we study to what extent do LLMs’ value profiles change in wanted ways (i.e. based on an instruction), while remaining robust to unwanted context-based change (i.e. based on different conversation topics). We instruct LLMs to simulate two populations: fictional characters and well-known real-world personas. In addition, we also study the coherence and robustness of LLMs’ value expression when they are not instructed to simulate any specific persona, corresponding to frequent real-world use cases. To our knowledge, this is the first study on value stability in LLMs.
@@
We use the Schwartz’s theory of Basic Personal Values (?, ?) and the corresponding Portrait Values Questionnaire (PVQ-40) (?, ?). This theory and questionnaire has been thoroughly studied and validated in the field of psychology, and it outlines ten basic personal values (Universalism, Benevolence, Conformity, Tradition, Security, Power, Achievement, Hedonism, Stimulation, and Self-Direction). Following that research, we outline two types of value stability studies in the psychology literature: Rank-Order (on a population/interpersonal level) and Ipsative (on an individual/intrapersonal level), see details below and on figures 2 and 3.
@@
  • Introduction and adaptation of the methodology for evaluating Rank-Order and Ipsative stability in LLMs
  • First analysis of value stability across contexts (including various conversation lengths)
  • Systematic comparison of Rank-Order and Ipsative stability of 19 LLMs with and without instructing the models to simulate specific personas
  • Analysis of stability in a downstream behavioral task
Figure 2: Rank-Order stability An example of estimating Rank-Order stability for benevolence. In each context, characters are ordered according to their benevolence scores in that context. In this example, the orders are almost the same in contexts 1 and 2 (high Rank-Order stability), and very different in contexts 2 and 3 (low Rank-Order stability).
Figure 3: Ipsative stability An example of estimating Ipsative stability for a character (Gandalf). Values are ordered according to the character’s scores in each context. In this example, the orders are the same in contexts 1 and 2 (high Ipsative stability), and different in contexts 2 and 3 (low Ipsative stability).
@@ 2 Methods
This section discusses how we administer the PVQ questionnaire over different contexts to evaluate value stability. We conduct experiments in two ways: with and without instructing the models to simulate specific personas. Different contexts are induced by simulating conversations on different topics with a separate instance of the same model (the interlocutor). To set a conversation topic (e.g. joke) we manually set the first interlocutor’s message (e.g. "Tell me a joke."). We let the models exchange $n$ messages, manually set the last interlocutor’s message as the query (e.g. PVQ item), and record the model’s response. After the questionnaire was given in each context, we estimate the Ipsative and Rank-Order stability. This process is shown in figure 1.
@@ 2.1 Administering the questionnaire
Administering the questionnaire consists of the following steps depicted in figure 1:
1. A model is instructed to simulate a persona (optional). We study personas from two populations: 1) 60 fictional characters from J.R.R. Tolkien’s universe , and 2) 50 real world personas (see appendix A for details). A persona setting instruction (e.g. “You are Gandalf from J. R. R. Tolkien’s Middle-earth legendarium.”) is given to the model (see appendix A for details).
2. A separate interlocutor model instance is created. The interlocutor model is an instance of the same model as the one being tested. The interlocutor is given the following instruction: “You are simulating a human using a chatbot. Your every reply must be in one sentence only.” If a persona was provided in step 1, the following sentence is added as the second sentence: “The chatbot is pretending to be character_name.”
3. A conversation topic is induced. The first interlocutor’s message is manually set to induce one of the following topics: grammar, joke, poem, history, chess. For example, to induce the topic of “joke” it is set to “Tell me a joke.”. See appendix A for details.
4. A conversation is simulated. The two models are let to exchange $n$ messages. In our experiments, $n$ is set to 3 (except when studying the effect of $n$ on stability).
5. A questionnaire is given. A questionnaire item is manually set as the interlocutor’s last message with a random order of suggested answers, and the model’s response is recorded. This is repeated for each question in parallel (with the same simulated conversation). That way, the model’s response is not influenced by responses to other questions.
6. A questionnaire is scored. The responses are scored to obtain the scores for the ten values. See appendix A for details.
7. Stability is estimated. If a persona was provided in step 1, steps 1 to 6 are repeated for every persona in the simulated population. Then, the whole process is repeated with five random seeds. Stability is estimated for each seed and then averaged, i.e. value stability for one model is estimated from 50/60k queries, depending on the population (5 seeds x 5 topics x 50/60 personas x 40 PVQ items). For reference, MMLU (?, ?) (a commonly used general knowledge benchmark) contains 14k test questions.
@@ 2.2 Estimating the stability
We estimate two types of value stability: Rank-Order and Ipsative. Rank-Order estimates the stability of some value at the population (interpersonal) level, as the stability of the order of participants in expressing that value. Intuitively, this can be seen as addressing the following question: "Does Jack always value Tradition more than Jane does?". Ipsative stability estimates the stability at the individual (intrapersonal) level as the stability of individuals value hierarchies. Intuitively, this can be seen as addressing the following question: "Does Jack always value Tradition more than Benevolence?".
@@
Rank-Order stability is used to estimate the stability of some value inside a population. In psychology, Rank-Order stability for some value can be computed as the correlation in the order of individuals’ scores at two points in time (Spearman correlation between the participants’ ranks). Here, instead of comparing the participant ranks at two points in time, we compare it in different contexts (see Fig. 2). We evaluate a model in five different contexts and compute the stability for each pair of contexts, and estimate the final stability as the average of those pairs.
@@
Ipsative stability is used to estimate the stability of an individual’s value profile. In psychology, Ipsative stability can be computed as the correlation between the ranks of values for the same individual at two points in time (Spearman correlation between the values’ scores). Here, instead of evaluating the value profile at two points in time, we evaluate it in different contexts (see Fig 3).
We evaluate models in five different contexts and compute the stability between each pair of contexts. We estimate the final stability by averaging over those pairs.
@@ 3 Experiments
This section provides an analysis of Ipsative and Rank-Order stability in LLMs. LLMs will be evaluated in two ways: with and without instructing the models to simulate particular personas. We aim to address the following questions:
@@
  • How does the stability of values expressed by LLMs compare to stability observed in human development?
@@
  • To what extent do conclusions made with PVQ transfer to a downstream behavioral task?

### v4

@@ Abstract
The standard way to study Large Language Models (LLMs) with benchmarks or psychology questionnaires is to provide many different queries from similar minimal contexts (e.g. multiple choice questions). However, due to LLMs’ highly context-dependent nature, conclusions from such minimal-context evaluations may be little informative about the model’s behavior in deployment (where it will be exposed to many new contexts). We argue that context-dependence (specifically, value stability) should be studied as a specific property of LLMs and used as another dimension of LLM comparison (alongside others such as cognitive abilities, knowledge, or model size). We present a case-study on the stability of value expression over different contexts (simulated conversations on different topics) as measured using a standard psychology questionnaire (PVQ) and on behavioral downstream tasks. Reusing methods from psychology, we study Rank-order stability on the population (interpersonal) level, and Ipsative stability on the individual (intrapersonal) level. We consider two settings (with and without instructing LLMs to simulate particular personas), two simulated populations, and three downstream tasks. We observe consistent trends in the stability of models and model families - Mixtral, Mistral, GPT-3.5 and Qwen families are more stable than LLaMa-2 and Phi. The consistency of these trends implies that some models exhibit higher value-stability than others, and that value stability can be estimated with the set of introduced methodological tools. When instructed to simulate particular personas, LLMs exhibit low Rank-order stability, which further diminishes with conversation length. This highlights the need for future research on LLMs that coherently simulate different personas. This paper provides a foundational step in that direction, and, to our knowledge, it is the first study of value stability in LLMs. The project website with code is available at https://sites.google.com/view/llmvaluestability.
Keywords: Large Language Models; LLM Stability; Value-stability, Basic Personal Values, Role-play, Context-dependence; Simulated populations
@@ 1 Introduction
In recent years, there has been an emergence of research using psychological tools to study Large Language Models (LLMs). In those studies, LLMs have often been used to simulate populations by instructing them to simulate different personas [1]. LLMs have also been used without providing such instructions, i.e. treated as a participant in a human study [2]. Questions in such studies have revolved around how Language Models express values [3], personality traits [4, 5], cognitive abilities [6], and how they could replicate human data [7].
@@
Figure 1: Evaluating the expressed value profile in context The tested LLM is prompted to play a specific role (e.g. Gandalf). We simulate a conversation on a topic (e.g. joke) with an interlocutor model (same LLM prompted to simulate a human user). Then, the tested LLM is given a psychology questionnaire (PVQ-40) aimed to assess its expressed values. We study the stability of these expressed values (and of behavior on downstream tasks) across diverse conversation topics and lengths. We consider the simulation of various fictional and real-world personas, as well as the case when the LLM is not prompted to play any particular persona. The messages and instructions in gray are set manually, and the messages in white are generated.
@@
The use of psychological tools with LLMs opens up many scientific questions, for instance regarding the nature of how the LLM-generated text depends on context, i.e. information present in the prompt or prior interaction with the model. Prompts or prior interaction, which we denote here as ’context’, can include any textual information such as instructions (e.g. personas to simulate), the dialogue history, stories written in specific styles, etc. Such contexts guide the generation of text by LLMs: different contexts may result in the expression of different behavior and values. This is sometimes beneficial and expected, e.g. an instruction to simulate some persona should influence the behavior and expressed values to be more aligned with that persona. However, this can also be problematic, e.g. a specific conversation topic may drastically influence the expressed behavior and values in unexpected ways, as we will highlight in experiments below. It should be noted that, depending on the application, different types of context-dependence can be beneficial or not. For example, to avoid malicious misuse of an AI-assistant, it might be beneficial to not express different values even when instructed to simulate different personas. On the other hand, simulating diverse populations of agents [8] or humans [7, 1] requires exactly this kind of context-dependence.
@@
The issue of unexpected context dependence in LLM is of crucial importance. Standard evaluation benchmarks, including those using psychological questionnaires to assess properties of LLMs, consist of sets of queries, often presented with a similar minimal context (e.g. knowledge or value-related questions presented as multiple choice questions with limited context). When deployed, LLMs are exposed to many new unforeseen contexts. This means that the standard benchmarks, by themselves, cannot estimate a model’s behavior in deployment (due to the LLMs’ highly context dependent nature). It is therefore crucial to evaluate the robustness of different models to unexpected context-based changes as a property, which can then be used as a dimension of LLM comparison alongside others such as cognitive abilities, knowledge, or model size.
@@
This challenge is particularly acute with the use of psychological questionnaires aimed at measuring psychological dimensions like values. Those tools were initially designed to probe humans, and thus make various assumptions about humans: for example, it is expected that the answers of most humans to questionnaires about value preferences should not be drastically influenced by the content of a random conversation. As we will show below, such an assumption does not hold for many LLMs, and thus strongly limits general interpretability of using these questionnaires in a context-independent manner. It is thus key to understand better how LLMs’ behavior (e.g. expression of values) may maintain coherence or change as a function of various kinds of contexts (ranging from explicit instruction to play a particular persona to discussions about topics that seem unrelated to the expressed psychological dimensions one studies).
@@
Previous research included certain experiments regarding unwanted context-based change (usually regarding syntactic changes in the prompt). These experiments led to conflicting results, sometimes showing robustness [9, 10, 4, 11], and sometimes sensitivity [2, 11]. These inconclusive results motivate research into the nature and the extent of context-dependence of various LLMs.
@@
In this paper, we present a case-study focusing on studying the stability of value expressed in 21 LLMs from 6 families. We simulate conversations with another instance of the LLM, which was instructed to simulate a human user (see Fig. 1). We study to what extent can LLMs simulate various personas in coherent ways, i.e. expressed values should change according to the instructed persona, but not based on the topic of a conversation not related to values. In other words, we study to what extent do LLM-expressed value profiles change in wanted ways (i.e. based on an instruction), while remaining robust to unwanted context-based change (i.e. based on different conversation topics). We instruct LLMs to simulate various fictional characters and real-world personas (from different countries and cultural backgrounds). We also study the coherence and robustness of LLMs’ value expression when they are not instructed to simulate any specific persona, corresponding to frequent real-world use cases. To our knowledge, this is the first study on value stability in LLMs.
@@
We use the Schwartz’s theory of Basic Personal Values [12, 13] and the corresponding Portrait Values Questionnaire (PVQ-40) [14]. This theory and questionnaire has been thoroughly studied and validated in the field of psychology, and it outlines universal ten basic personal values (Universalism, Benevolence, Conformity, Tradition, Security, Power, Achievement, Hedonism, Stimulation, and Self-Direction). The PVQ-40 questionnaire has been used to study values in a large diversity of countries and cultures [15]. Refer to Appendix A for details. We outline two types of value stability commonly studied in the field of psychology: Rank-order (on a population/interpersonal level) and Ipsative (on an individual/intrapersonal level) (see Figures 2 and 3).
Figure 2: Rank-order stability An example of estimating Rank-order stability of benevolence. In each context, characters are ordered according to their benevolence scores in that context. In this example, the orders are almost the same in contexts 1 and 2 (high Rank-order stability), and very different in contexts 2 and 3 (low Rank-order stability).
Figure 3: Ipsative stability An example of estimating Ipsative stability for a fictional character (Gandalf). Values are ordered according to the character’s scores in each context. In this example, the orders are the same in contexts 1 and 2 (high Ipsative stability), and different in contexts 2 and 3 (low Ipsative stability).
@@
  • Conceptualization of context-dependence (specifically, value-stability) as a property of LLM, to be used as another dimension of LLM comparison alongside others such as cognitive abilities, model size or inference speed.
  • A set of methodological tools enabling the study of (intrapersonal and interpersonal) value stability and context-dependence in LLMs
  • First analysis of stability of value expression in LLMs across contexts. The analysis was conducted over: 21 LLMs from six families, two types of stability, two settings (with and without instructing them to simulate specific personas), two simulated populations, and three downstream tasks.
2 Related Work
There has been a growing number of works using psychological questionnaires to study LLMs. [2] evaluated GPT-3 on a battery of vignette-based cognitive tests, and [6] evaluated a LaMDa model on a battery of classical developmental tasks and compared its performance to that of human children. Some works have evaluated LLMs using personality questionnaires [5, 16] and tests for creativity [17] Although not directly using psychological questionnaires, there are works heavily inspired by psychology which estimate LLMs’ Theory of Mind through textual representations of standard False-belief tasks [18, 19]. Likewise, psychological questionnaires have been used with LLMs to inform psychological [20] and neuroscience [21, 22] hypotheses.
A body of work has studied LMs as simulating a diversity of individuals and cultures [23]. [24] train LSTM-based [25] conversational models to simulate different personas. [26] introduced the metaphor of role-playing, where an LLM always chooses a character to role-play based on context. Cultural expression has been studied by inducing personalities from different countries [27] and by querying the model in different languages [28]. [29] also used those methods to compare LLM responses to human data from the World Values Survey [30] and the Pew Global Attitudes Survey [31]. [10] induced personas of various demographic groups and observed a left-leaning bias. [32] fine-tune LLM to simulate different high- and low-resource cultures. [33] study the cultural evolution of stories in populations of LLMs induced with different personalities. [34] simulate children of different ages to recreate developmental stages of exploration, and simulate different domain experts to improve question answering. [35] show that inducing famous personalities crease toxicity. [7] replicate psychological studies with humans by varying the participants’ names to simulate a culturally diverse population. Similarly, [1] replicate data from human studies by prompting the model with backstories of real human participants in those original studies. LLMs have also been used to simulate students in order to train teachers [36].
Previous work has also highlighted the influence of context on LLMs behavior [37]. [38] show that exposing an LLM to some statement increases its perceived truthfulness at a later time, and [39] demonstrate the tendency of models to repeat back the user’s answer. [40] highlights the need for multi-turn evaluation to uncover attractors in the cultural evolution of stories generated by LLMs. The most similar work to ours is a concurrent paper studying the coherence of simulated personas in general (as opposed to our specific focus on personal value stability in simulated individuals and populations). That work proposes to increase the similarity of LLaMa-2-70b-chat [41] model’s answers before and after simulated conversations by reweighting the instruction’s attention weights [42].
@@ 3 Methods
This section discusses how we administer the PVQ questionnaire over different contexts to evaluate value stability. We conduct experiments in two ways: with and without instructing the models to simulate specific personas. Different contexts are induced by simulating conversations on different topics with a separate instance of the same model (the interlocutor). To set a conversation topic (e.g. joke) we manually set the first interlocutor’s message (e.g. "Tell me a joke."). We let the models exchange $n$ messages, manually set the last interlocutor’s message as the query (e.g. PVQ item), and record the model’s response. After the questionnaire was given in each context, we estimate the Ipsative and Rank-order stability. This process is shown on Fig 1.
@@ 3.1 Administering the questionnaire
Administering the questionnaire consists of the following steps depicted in Fig 1:
1. A model is instructed to simulate a persona (optional). We study personas from two populations: 1) 60 fictional characters from J.R.R. Tolkien’s universe , and 2) 50 real-world personas (see Appendix B.1 for details). Instructing a model to simulate a well-known persona is a simple way to precisely define a persona’s character (assuming the models were trained on enough of related data). A persona setting instruction (e.g. “You are Gandalf from J. R. R. Tolkien’s Middle-earth legendarium.”) is given to the model (see Appendix B.2 for details).
2. A separate interlocutor model instance is created. The interlocutor model is an instance of the same model as the one being tested. The interlocutor is given the following instruction: “You are simulating a human using a chatbot. Your every reply must be in one sentence only.” If a persona was provided in step 1, the following sentence is added as the second sentence: “The chatbot is pretending to be <character_name>.”
3. A conversation topic is induced. The first interlocutor’s message is manually set to induce one of the following topics: grammar, joke, poem, history, chess. For example, to induce the topic of “joke” it is set to “Tell me a joke.”. See Appendix B.3 for details.
4. A conversation is simulated. The two models are let to exchange $n$ messages.  [^1] The default parameters from each model’s card on the huggingface website (https://huggingface.co/models) are used. In our experiments, $n$ is set to 3 (except when studying the effect of $n$ on stability).
5. A questionnaire is given. A questionnaire item is manually set as the interlocutor’s last message (with a random order of suggested answers), and the model’s response is recorded. This is repeated for each question in parallel (with the same simulated conversation). That way, the model’s response is not influenced by the responses to other questions.
6. A questionnaire is scored. The responses are scored to obtain the scores for the ten values. See Appendix B.5 for details.
7. Stability is estimated. If a persona was provided in step 1, steps 1 to 6 are repeated for every persona in the simulated population. Then, the whole process is repeated with five random seeds. Stability is estimated for each seed and then averaged, i.e. value stability for one model is estimated from 50/60k queries, depending on the population (5 seeds x 5 topics x 50/60 personas x 40 PVQ items). For reference, MMLU [43] (a commonly used general knowledge benchmark) contains 14k test questions.
@@ 3.2 Estimating the stability
We estimate two types of value stability: Rank-order and Ipsative. Rank-order estimates the stability of some value at the population (interpersonal) level, and Ipsative stability estimates the stability of value profiles at the individual (intrapersonal) level.
@@
Rank-order stability is used to estimate the stability of some value inside a population. In psychology, it is computed as the correlation in the order of individuals at two points in time (individuals are ordered based on their expression of that value). Intuitively, this can be seen as addressing the following question: "Does Jack always value Tradition more than Jane does?". As shown in Fig 2, instead of comparing two points in time, we compare different contexts (simulated conversations of different topics).
@@
Ipsative stability is used to estimate the stability of an individual’s value profile. In psychology, it is computed as the correlation in the order of values expressed by the same individual at two points in time. Intuitively, this can be seen as addressing the following question: "Does Jack always value Tradition more than Benevolence?". As shown in Fig 3, instead of evaluating the value profile at two points in time, we evaluate it in different contexts (simulated conversations of different topics).
For both Rank-order and Ipsative stability, we evaluate models in five different contexts, compute the stability between each pair of contexts, and estimate the final stability by averaging over those pairs. For Rank-order stability, we additionally average the stabilities of the ten values.
@@ 4 Experiments
This section provides an analysis of context-dependence (specifically, value-stability) as a property of LLMs. We observe consistent trends along different experimental settings, implying that some models and model families exhibit more favorable context-dependence compared to others. LLMs are evaluated in two ways: with and without instructing them to simulate particular personas. We aim to address the following questions:
@@
  • How does the stability of values expressed by LLMs compare to that observed in human development?
@@
  • To what extent do conclusions made with PVQ transfer to downstream behavioral tasks?
  • Is value expression on PVQ correlated with behavior on a downstream task?
  • How additional contexts affect stability estimates?